I"m stuck here 4 2/3 - 3 3/5
Answer by checkley79(3341)

4 2/3 - 3 3/5
convert to improper fractions
4 2/3 = 14/3
3 3/5 = 18/5
14/3-18/5
find a common denominator
3*5=15
14/3=5*14/15 = 70/15
18/5=18*3/15 = 54/15
now subtract
(70-54)/15 =16/15 ans.